# Furstenberg set problem and exceptional set estimate in prime fields: dimension two implies higher dimensions

### v1

@@ 4. Furstenberg set problem: Lower bound I
#$\#E\gtrsim_{\varepsilon}\sum_{\xi\in\Xi_{\circ}}p^{\mathbf{F}(s_{2},t_{\xi};k,k-1)-\varepsilon}\geq\#\Xi_{\circ}\cdot p^{\mathbf{F}(s_{2},t_{\circ};k,k-1)-\varepsilon}\gtrapprox p^{u}p^{\mathbf{F}(s_{2},t_{\circ};k,k-1)+t_{1}+v-t_{\circ}-\varepsilon}.$
By Lemma 20, we have
#$\#E\gtrapprox_{\varepsilon}p^{u+\mathbf{F}(s_{2},t_{1}+v;k,k-1)-\varepsilon}.$
The next lemma is a recursive inequality of $\mathbf{F}(s,t;k+1,k)$, which can be used to bound the right hand side of (31).
Lemma 32.
@@
(38) $\begin{cases}t_{1}\in[0,k-1],t_{2}\in[0,2];\\ s_{1}\in[0,1],s_{2}\in[0,s];\\ u\in[s_{1},1],v\in[0,1];\end{cases}$
@@
(39) $\begin{cases}t_{1}+t_{2}\geq t;\\ s_{1}+s_{2}\geq s;\\ u+v\geq\mathbf{F}(s_{1},t_{2};2,1);\end{cases}$
@@
(40) $u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}\geq\mathbf{F}(s,t;k+1,k).$
Let us quickly see how Lemma 30 and Lemma 32 combined to prove Proposition 29.
Proof of Proposition 29 assuming Lemma 30 and Lemma 32.
If $s=0$ or $t=0$, by Proposition 27 or 28, we are done. So, we suppose $s,t>0$.
@@
(41) #$\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;k+1,k)-C\varepsilon}$
@@
Suppose (27) holds. Then in particular, Conjecture 4 also holds. If $E$ is a $(s,t;k+1,k)$-set, by Lemma 30, there exist numbers $t_{1},t_{2},s_{1},s_{2},u,v$, such that
(42) $\begin{cases}t_{1}\in[0,k-1],t_{2}\in[0,2];\\ s_{1}\in[0,1],s_{2}\in[0,s];\\ u\in[s_{1}-\varepsilon,1],v\in[0,1];\end{cases}$
@@
(43) $\begin{cases}t_{1}+t_{2}\geq t-\varepsilon;\\ s_{1}+s_{2}\geq s-\varepsilon;\\ u+v\geq\mathbf{F}(s_{1},t_{2};2,1)-\varepsilon;\end{cases}$
@@
(44) #$\#E\gtrsim_{\varepsilon}p^{u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}-\varepsilon}.$
Here, we used the fact that when $p$ is large enough (depending on $\varepsilon$), then $p^{t_{1}+t_{2}}\gtrapprox p^{t}$ implies $t_{1}+t_{2}\geq t-\varepsilon$. Also, by our assumption. $t-\varepsilon,s-\varepsilon>0$. Let $t^{\prime}=t-\varepsilon$. $s^{\prime}=s-\varepsilon$.
@@
(45) $\begin{cases}t_{1}\in[0,k-1],t_{2}\in[0,2];\\ s_{1}\in[0,1],s_{2}\in[0,s];\\ u^{\prime}\in[s_{1},1],v^{\prime}\in[0,1];\end{cases}$
@@
(46) $\begin{cases}t_{1}+t_{2}\geq t^{\prime};\\ s_{1}+s_{2}\geq s^{\prime};\\ u^{\prime}+v^{\prime}\geq\mathbf{F}(s_{1},t_{2};2,1);\end{cases}$
@@
(47) #$\#E\gtrsim p^{u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}-\varepsilon}\geq p^{u^{\prime}+\max\{\mathbf{F}(s_{2},t_{1}+v^{\prime};k,k-1),s_{2}+v^{\prime}\}-O(\varepsilon)}.$
In the last inequality, we used Lemma 20.
Finally, by Lemma 32, we have
@@
The last inequality is by Proposition 21 and 20.
@@
#$\#E\gtrsim p^{\mathbf{F}(s,t;k+1,k)-C\varepsilon}.$
It remains to prove Lemma 32. It can be reduced to the following lemma, with $\geq$” in (39) all replaced by “$=$”.
Lemma 33.
@@
(48) $\begin{cases}t_{1}\in[0,k-1],t_{2}\in[0,2];\\ s_{1}\in[0,1],s_{2}\in[0,s];\\ u\in[s_{1},1],v\in[0,1];\end{cases}$
@@
(49) $\begin{cases}t_{1}+t_{2}=t;\\ s_{1}+s_{2}=s;\\ u+v=\mathbf{F}(s_{1},t_{2};2,1);\end{cases}$
@@
(50) $u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}\geq\mathbf{F}(s,t;k+1,k).$
We show that Lemma 33 implies Lemma 32.
Proof of Lemma 32 assuming Lemma 33.
The idea is to slightly modify $s_{1},s_{2},t_{1},t_{2},u,v$ so that the $\geq$” in (39) become “$=$”, and then we can apply Lemma 33.
For $i=1,2$, we can find $t_{i}^{\prime}\in[0,t_{i}],s^{\prime}_{i}\in[0,s_{i}]$, so that $t_{1}^{\prime}+t_{2}^{\prime}=t,s_{1}^{\prime}+s_{2}^{\prime}=s$. It is slightly trickier to modify $u,v$. Noting that $u+v\geq\mathbf{F}(s_{1},t_{2};2,1)\geq\mathbf{F}(s_{1}^{\prime},t_{2}^{\prime};2,1)\geq s_{1}^{\prime}$, we can find $u^{\prime}\in[s_{1}^{\prime},u],v^{\prime}\in[0,v]$ so that $u^{\prime}+v^{\prime}=\mathbf{F}(s_{1}^{\prime},t_{2}^{\prime};2,1)$. Now we can apply Lemma 33 to the numbers $t_{1}^{\prime},t_{2}^{\prime},s_{1}^{\prime},s_{2}^{\prime},u^{\prime},v^{\prime}$. We obtain
@@
It remains to prove Lemma 33.
Proof of Lemma 33.
@@
If $t_{2}\geq\tau$, then $\mathbf{F}(\sigma,t_{2};2,1)\geq\mathbf{F}(\sigma,\tau;2,1)$. If $t_{2}\leq\tau$, then $\mathbf{F}(\sigma,t_{2};2,1)+\tau-t_{2}\geq\mathbf{F}(\sigma,\tau;2,1)$, by Lemma 20. Therefore, we have
@@
(51) $\begin{split}\mathbf{F}(s_{2},t_{1}+v;k,k-1)=\mathbf{F}(d+(\sigma-s_{1}),t_{1}+v;k,k-1)\\ =d+\min\{\sigma-s_{1}+\tau^{\prime},\frac{3}{2}(\sigma-s_{1})+\frac{1}{2}\tau^{\prime},(\sigma-s_{1})+1\}.\end{split}$
@@
(52) $\begin{split}\mathbf{F}&\geq u+\max\{s_{2}+\tau^{\prime},s_{2}+v\}\\ &=u+s_{2}+\max\{\tau^{\prime},v\}.\end{split}$
@@
5. Furstenberg set problem: Lower bound II
Proposition 34.
@@ 5. Furstenberg set problem: Lower bound II
(53) $\inf_{E:\ (s,t;n^{\prime},k)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n^{\prime},k)-\varepsilon},$
@@
(54) $\inf_{E:\ (s,t;n,k)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-\varepsilon},$
@@
We prove the following lemma.
Lemma 35.
@@
(55) $\begin{cases}1\leq p^{t_{1}}\leq p^{(k+1)(n-k-1)},1\leq p^{t_{2}}\leq p^{k+1};\\ p^{s}\lesssim p^{s_{1}}\leq p^{k},1\leq p^{s_{2}}\leq p;\end{cases}$
@@
(56) $\begin{cases}p^{t_{1}+t_{2}}\gtrapprox p^{t};\\ p^{s_{1}+s_{2}}\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t_{2};n-1,k)-\varepsilon};\end{cases}$
@@
(57) #$\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{1};n-1,k)+s_{2}-\varepsilon}.$
@@
where $\mathcal{V}\subset A(k,\mathbb{F}_{p}^{n})$ with #$\#\mathcal{V}\sim p^{t}$, and $Y(V)\subset V$ with #$\#Y(V)\sim p^{s}$.
@@
By properly choosing the coordinate, we can assume $\gtrsim 1$ fraction of the $k$-planes in $\mathcal{V}$ are not parallel to the $n$-th axis $\mathbf{0}^{n-1}\times\mathbb{F}_{p}$. Throwing away some $V\in\mathcal{V}$, we can just assume all the $k$-planes in $\mathcal{V}$ are not parallel to $\mathbf{0}^{n-1}\times\mathbb{F}_{p}$ with $\mathcal{V}$ still satisfying #$\#\mathcal{V}\gtrsim p^{t}$. Because of this, we see that for any $V\in\mathcal{V}$, the projection of $V$ onto $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$, denoted by $\pi_{\mathbb{F}_{p}^{n-1}}(V)$ is a $k$-plane in $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$.
@@
First pigeonholing. We assume there exists $\mathcal{W}\subset A(k,\mathbb{F}_{p}^{n-1}\times\mathbf{0})$ with #$\#\mathcal{W}\sim p^{t_{1}}$ such that for each $W\in\mathcal{W}$, #$\#\mathcal{V}_{W}\sim p^{t_{2}}$. Here, $1\leq p^{t_{1}}\leq p^{(k+1)(n-k-1)},1\leq p^{t_{2}}\leq p^{k+1}$ are dyadic numbers and
(58) $p^{t_{1}}p^{t_{2}}\gtrapprox p^{t}.$
The upper bound for $t_{1},t_{2}$ is because #$\#A(k,\mathbb{F}_{p}^{n-1}\times\mathbf{0})\lesssim p^{(k+1)(n-k-1)}$, and #$\#\mathcal{V}_{W}\leq\#A(k,W\times\mathbb{F}_{p})\lesssim p^{k+1}$. We still use $\mathcal{V}$ to denote $\bigcup_{W\in\mathcal{W}}\mathcal{V}_{W}$. And we see that
#$\#\mathcal{V}\gtrapprox p^{t}.$
Fix a $W\in\mathcal{W}$. We want to estimate $\bigcup_{V\in\mathcal{V}_{W}}Y(V)(\subset W\times\mathbb{F}_{p}\cong\mathbb{F}_{p}^{k+1})$. This is some sort of Furstenberg set for $k$-planes in $(k+1)$-dimensional space. Also, an important preperty is that each $V\in\mathcal{V}_{W}$ is not parallel to the last coordinate. Therefore, any line parallel to $\mathbf{0}^{n-1}\times\mathbb{F}_{p}$ will intersect $V$ at a single point. We will exploit some “quasi-product” structure of this set.
Lemma 36.
Assume (53) holds for $n^{\prime}=k+1$. Suppose $E=\bigcup_{V\in\mathcal{V}}Y(V)$ is a $(s,t_{2};k,k+1)$-set. Also, we assume each $V\in\mathcal{V}$ intersects $\mathbf{0}^{k}\times\mathbb{F}_{p}$ at a single point. Then there exist dyadic numbers $p^{s_{1}},p^{s_{2}}$ such that
@@
for any $\varepsilon>0$. Also, $E$ contains a subset of form
@@
where $\Xi\subset\mathbb{F}_{p}^{k}\times\mathbf{0}$ with #$\#\Xi\sim p^{s_{1}}$, and $Y_{\xi}\subset E\cap\bigg{(}\xi+(\mathbf{0}^{k}\times\mathbb{F}_{p})\bigg{)}$ with #$\#Y_{\xi}\sim p^{s_{2}}$.
@@
For $\xi\in\mathbb{F}_{p}^{k}\times\mathbf{0}$, define $E(\xi):=E\cap\bigg{(}\xi+(\mathbf{0}^{k}\times\mathbb{F}_{p})\bigg{)}$. By pigeonholing on vertical slices of $E$, we can find $\Xi\subset\mathbb{F}_{p}^{k}\times\mathbf{0}$ so that for all $\xi\in\Xi$, $E(\xi)$ have comparable cardinality. We may denote #$\#\Xi\sim p^{s_{1}},\#E(\xi)\sim p^{s_{2}}$ for two dyadic numbers $p^{s_{1}},p^{s_{2}}\geq 1$. We can also assume $p^{s_{1}}p^{s_{2}}\gtrapprox\#E$. Hence, assuming (53) holds for $n^{\prime}=k+1$, we have
@@
However, we still need the condition $p^{s_{1}}\gtrsim p^{s}$. This is handled by the same trick as in the proof of Lemma 31. We will conduct an algorithm to find a sequence $E=E_{1}\supset E_{2}\supset\cdots\supset E_{m}$, and disjoint subsets $\Xi_{1},\Xi_{2}\cdots,\Xi_{m}$ of $\mathbb{F}_{p}^{k}\times\mathbf{0}$. To start with, let $E_{1}=E$. By pigeonholing, we can find $\Xi_{1}\subset\mathbb{F}_{p}^{k}\times\mathbf{0}$, such that #$\#\Xi_{1}\sim p^{u_{1}}$, #$\#E(\xi)\sim p^{v_{1}}$ for each $\xi\in\Xi$, and
@@
Suppose we have defined $E_{i},\Xi_{i}$ for $11\leq i\leq j-$ 1. If $\sum_{i=1}^{j-1}p^{u_{i}}\ll p^{s}$, then we define
@@
Since each $V\in\mathcal{V}$ intersect $\xi+(\mathbf{0}^{k}\times\mathbb{F}_{p})$ (and hence $E(\xi)$) at a single point, we see that $E_{j}$ is still a $(s,t_{2};2,1)$-set. By the same pigeonholing, we find $\Xi_{j}\subset(\mathbb{F}_{p}^{k}\times\mathbf{0})\setminus\bigcup_{1\leq i\leq j-1}\Xi_{i}$, so that #$\#\Xi_{j}\sim p^{u_{j}}$, #$\#E(\xi)\sim p^{v_{j}}$ for each $\xi\in\Xi_{j}$, and
@@
If $\sum_{i=1}^{j-1}p^{u_{i}}\gtrsim p^{s}$, then we stop and let $m=j-1$.
The algorithm will stop in finite steps. Denote $\Xi:=\bigcup_{1\leq i\leq m}\Xi_{i}$. Let $p^{s_{1}},p^{s_{2}}$ be dyadic numbers so that $p^{s_{2}}=\min_{1\leq i\leq m}p^{v_{i}}$, $p^{s_{1}}\sim\#\Xi$. For each $\xi\in\Xi$, let $Y_{\xi}$ be a subset of $E(\xi)$ with #$\#Y_{\xi}\sim p^{s_{2}}$. This is doable, since #$\#E(\xi)\gtrsim p^{s_{2}}$ if $\xi\in\Xi$. One can check

### v2

@@ 4 Furstenberg set problem: Lower bound I
$E\gtrsim_{\varepsilon}\sum_{\xi\in\Xi_{\circ}}p^{\mathbf{F}(s_{2},t_{\xi};k,k-1)-\varepsilon}\geq\#\Xi_{\circ}\cdot p^{\mathbf{F}(s_{2},t_{\circ};k,k-1)-\varepsilon}\gtrapprox p^{u}p^{\mathbf{F}(s_{2},t_{\circ};k,k-1)+t_{1}+v-t_{\circ}-\varepsilon}.$
By Lemma 2.11, we have
$E\gtrapprox_{\varepsilon}p^{u+\mathbf{F}(s_{2},t_{1}+v;k,k-1)-\varepsilon}.$
Note that the implicit constants in the above arguments may depend on various parameters $s_{1},s_{2},t_{1},t_{2}$ and so on. To get rid of this dependence, we use the trick discussed before this lemma. Hence, we update the estimate to
$p^{u+v}\gtrsim_{\varepsilon}p^{\mathbf{F}((s_{1}-\varepsilon^{2})_{+},(t_{2}-\varepsilon^{2})_{+};2,1)-\varepsilon},$
$E\gtrsim_{\varepsilon}p^{u+\max\{\mathbf{F}((s_{2}-\varepsilon^{2})_{+},(t_{1}+v-\varepsilon^{2})_{+};k,k-1),s_{2}+v\}-\varepsilon}.$
Now, the implicit constants only depend on $s,t,k,\varepsilon$. ∎
The next lemma is a recursive inequality of $\mathbf{F}(s,t;k+1,k)$, which can be used to bound the right-hand side of (31).
Lemma 4.6.
@@
$\begin{cases}t_{1}\in[0,k-1],t_{2}\in[0,2];\\ s_{1}\in[0,1],s_{2}\in[0,s];\\ u\in[s_{1},1],v\in[0,1];\end{cases}$ (38)
@@
$\begin{cases}t_{1}+t_{2}\geq t;\\ s_{1}+s_{2}\geq s;\\ u+v\geq\mathbf{F}(s_{1},t_{2};2,1);\end{cases}$ (39)
@@
$u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}\geq\mathbf{F}(s,t;k+1,k).$ (40)
Let us quickly see how Lemma 4.4 and Lemma 4.6 combined to prove Proposition 4.3.
Proof of Proposition 4.3 assuming Lemma 4.4 and Lemma 4.6.
If $s=0$ or $t=0$, by Proposition 4.1 or 4.2, we are done. So, we suppose $s,t>0$.
@@
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;k+1,k)-C\varepsilon}$ (41)
@@
Suppose (27) holds. Then in particular, Conjecture 1.5 also holds. If $E$ is a $(s,t;k+1,k)$-set, by Lemma 4.4, there exist numbers $t_{1},t_{2},s_{1},s_{2},u,v$, such that
$\begin{cases}t_{1}\in[0,k-1],t_{2}\in[0,2];\\ s_{1}\in[0,1],s_{2}\in[0,s];\\ u\in[s_{1}-\varepsilon,1],v\in[0,1];\end{cases}$ (42)
@@
$\begin{cases}t_{1}+t_{2}\geq t-\varepsilon;\\ s_{1}+s_{2}\geq s-\varepsilon;\\ u+v\geq\mathbf{F}(s_{1},t_{2};2,1)-\varepsilon;\end{cases}$ (43)
@@
$E\gtrsim_{\varepsilon}p^{u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}-\varepsilon}.$ (44)
Here, we renamed $t_{1}$ to be $(t_{1}-\varepsilon^{2})_{+}$ in Lemma 4.4 (and similar for $t_{2},s_{1},s_{2},u,v$). Also, we used the fact that when $p$ is large enough (depending on $\varepsilon$), then $p^{t_{1}+t_{2}}\gtrapprox p^{t}$ implies $t_{1}+t_{2}\geq t-\varepsilon$. Also, by our assumption. $t-\varepsilon,s-\varepsilon>0$. Let $t^{\prime}=t-\varepsilon$. $s^{\prime}=s-\varepsilon$.
@@
$\begin{cases}t_{1}\in[0,k-1],t_{2}\in[0,2];\\ s_{1}\in[0,1],s_{2}\in[0,s];\\ u^{\prime}\in[s_{1},1],v^{\prime}\in[0,1];\end{cases}$ (45)
@@
$\begin{cases}t_{1}+t_{2}\geq t^{\prime};\\ s_{1}+s_{2}\geq s^{\prime};\\ u^{\prime}+v^{\prime}\geq\mathbf{F}(s_{1},t_{2};2,1);\end{cases}$ (46)
@@
$E\gtrsim p^{u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}-\varepsilon}\geq p^{u^{\prime}+\max\{\mathbf{F}(s_{2},t_{1}+v^{\prime};k,k-1),s_{2}+v^{\prime}\}-O(\varepsilon)}.$ (47)
In the last inequality, we used Lemma 2.11.
Finally, by Lemma 4.6, we have
@@
The last inequality is by Proposition 2.12 and 2.11.
@@
$E\gtrsim p^{\mathbf{F}(s,t;k+1,k)-C\varepsilon}.$
It remains to prove Lemma 4.6. It can be reduced to the following lemma, with $\geq$” in (39) all replaced by “==”.
Lemma 4.7.
@@
$\begin{cases}t_{1}\in[0,k-1],t_{2}\in[0,2];\\ s_{1}\in[0,1],s_{2}\in[0,s];\\ u\in[s_{1},1],v\in[0,1];\end{cases}$ (48)
@@
$\begin{cases}t_{1}+t_{2}=t;\\ s_{1}+s_{2}=s;\\ u+v=\mathbf{F}(s_{1},t_{2};2,1);\end{cases}$ (49)
@@
$u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}\geq\mathbf{F}(s,t;k+1,k).$ (50)
We show that Lemma 4.7 implies Lemma 4.6.
Proof of Lemma 4.6 assuming Lemma 4.7.
The idea is to slightly modify $s_{1},s_{2},t_{1},t_{2},u,v$ so that the $\geq$” in (39) becomes “==”, and then we can apply Lemma 4.7.
For $i=1,2$, we can find $t_{i}^{\prime}\in[0,t_{i}],s^{\prime}_{i}\in[0,s_{i}]$, so that $t_{1}^{\prime}+t_{2}^{\prime}=t,s_{1}^{\prime}+s_{2}^{\prime}=s$. It is slightly trickier to modify $u,v$. Noting that $u+v\geq\mathbf{F}(s_{1},t_{2};2,1)\geq\mathbf{F}(s_{1}^{\prime},t_{2}^{\prime};2,1)\geq s_{1}^{\prime}$, we can find $u^{\prime}\in[s_{1}^{\prime},u],v^{\prime}\in[0,v]$ so that $u^{\prime}+v^{\prime}=\mathbf{F}(s_{1}^{\prime},t_{2}^{\prime};2,1)$. Now we can apply Lemma 4.7 to the numbers $t_{1}^{\prime},t_{2}^{\prime},s_{1}^{\prime},s_{2}^{\prime},u^{\prime},v^{\prime}$. We obtain
@@
It remains to prove Lemma 4.7.
Proof of Lemma 4.7.
@@
If $t_{2}\geq\tau$, then $\mathbf{F}(\sigma,t_{2};2,1)\geq\mathbf{F}(\sigma,\tau;2,1)$. If $t_{2}\leq\tau$, then $\mathbf{F}(\sigma,t_{2};2,1)+\tau-t_{2}\geq\mathbf{F}(\sigma,\tau;2,1)$, by Lemma 2.11. Therefore, we have
@@
$\begin{split}\mathbf{F}(s_{2},t_{1}+v;k,k-1)=\mathbf{F}(d+(\sigma-s_{1}),t_{1}+v;k,k-1)\\ =d+\min\{\sigma-s_{1}+\tau^{\prime},\frac{3}{2}(\sigma-s_{1})+\frac{1}{2}\tau^{\prime},(\sigma-s_{1})+1\}.\end{split}$ (51)
@@
$\begin{split}\mathbf{F}&\geq u+\max\{s_{2}+\tau^{\prime},s_{2}+v\}\\ &=u+s_{2}+\max\{\tau^{\prime},v\}.\end{split}$ (52)
@@
5 Furstenberg set problem: Lower bound II
Proposition 5.1.
@@ 5 Furstenberg set problem: Lower bound II
$\inf_{E:\ (s,t;n^{\prime},k)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n^{\prime},k)-\varepsilon},$ (53)
@@
$\inf_{E:\ (s,t;n,k)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-\varepsilon},$ (54)
@@
We prove the following lemma. Actually, we also need to deal with a technical thing as mentioned before Lemma 4.4. Since we have done it once in the proof of Lemma 4.4, we will not reproduce here.
Lemma 5.2.
@@
$\begin{cases}1\leq p^{t_{1}}\leq p^{(k+1)(n-k-1)},1\leq p^{t_{2}}\leq p^{k+1};\\ p^{s}\lesssim p^{s_{1}}\leq p^{k},1\leq p^{s_{2}}\leq p;\end{cases}$ (55)
@@
$\begin{cases}p^{t_{1}+t_{2}}\gtrapprox p^{t};\\ p^{s_{1}+s_{2}}\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t_{2};n-1,k)-\varepsilon};\end{cases}$ (56)
@@
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{1};n-1,k)+s_{2}-\varepsilon}.$ (57)
@@
where $\mathcal{V}\subset A(k,\mathbb{F}_{p}^{n})$ with $\mathcal{V}\sim p^{t}$, and $Y(V)\subset V$ with $Y(V)\sim p^{s}$.
@@
By properly choosing the coordinate, we can assume that $\gtrsim 1$ fraction of the $k$-planes in $\mathcal{V}$ are not parallel to the $n$-th axis $\mathbf{0}^{n-1}\times\mathbb{F}_{p}$. Throwing away some $V\in\mathcal{V}$, we can just assume all the $k$-planes in $\mathcal{V}$ are not parallel to $\mathbf{0}^{n-1}\times\mathbb{F}_{p}$ with $\mathcal{V}$ still satisfying $\mathcal{V}\gtrsim p^{t}$. Because of this, we see that for any $V\in\mathcal{V}$, the projection of $V$ onto $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$, denoted by $\pi_{\mathbb{F}_{p}^{n-1}}(V)$ is a $k$-plane in $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$.
@@
First pigeonholing. We assume that there exists $\mathcal{W}\subset A(k,\mathbb{F}_{p}^{n-1}\times\mathbf{0})$ with $\mathcal{W}\sim p^{t_{1}}$ such that for each $W\in\mathcal{W}$, $\mathcal{V}_{W}\sim p^{t_{2}}$. Here, $1\leq p^{t_{1}}\leq p^{(k+1)(n-k-1)},1\leq p^{t_{2}}\leq p^{k+1}$ are dyadic numbers, and
$p^{t_{1}}p^{t_{2}}\gtrapprox p^{t}.$ (58)
The upper bound for $t_{1},t_{2}$ is because $A(k,\mathbb{F}_{p}^{n-1}\times\mathbf{0})\lesssim p^{(k+1)(n-k-1)}$, and $\mathcal{V}_{W}\leq\#A(k,W\times\mathbb{F}_{p})\lesssim p^{k+1}$. We still use $\mathcal{V}$ to denote $\bigcup_{W\in\mathcal{W}}\mathcal{V}_{W}$. And we see that
$\mathcal{V}\gtrapprox p^{t}.$
Fix a $W\in\mathcal{W}$. We want to estimate $\bigcup_{V\in\mathcal{V}_{W}}Y(V)(\subset W\times\mathbb{F}_{p}\cong\mathbb{F}_{p}^{k+1})$. This is some sort of Furstenberg set for $k$-planes in $(k+1)$-dimensional space. Also, an important property is that each $V\in\mathcal{V}_{W}$ is not parallel to the last coordinate. Therefore, any line parallel to $\mathbf{0}^{n-1}\times\mathbb{F}_{p}$ will intersect $V$ at a single point. We will exploit some “quasi-product” structure of this set.
Lemma 5.3.
Assume (53) holds for $n^{\prime}=k+1$. Suppose $E=\bigcup_{V\in\mathcal{V}}Y(V)$ is a $(s,t_{2};k,k+1)$-set. In addition, we assume that each $V\in\mathcal{V}$ intersects $\mathbf{0}^{k}\times\mathbb{F}_{p}$ at a single point. Then there exist dyadic numbers $p^{s_{1}},p^{s_{2}}$ such that
@@
for any $\varepsilon>0$. In addition, $E$ contains a subset of form
@@
where $\Xi\subset\mathbb{F}_{p}^{k}\times\mathbf{0}$ with $\Xi\sim p^{s_{1}}$, and $Y_{\xi}\subset E\cap\bigg(\xi+(\mathbf{0}^{k}\times\mathbb{F}_{p})\bigg)$ with $Y_{\xi}\sim p^{s_{2}}$.
@@
For $\xi\in\mathbb{F}_{p}^{k}\times\mathbf{0}$, define $E(\xi):=E\cap\bigg(\xi+(\mathbf{0}^{k}\times\mathbb{F}_{p})\bigg)$. By pigeonholing on vertical slices of $E$, we can find $\Xi\subset\mathbb{F}_{p}^{k}\times\mathbf{0}$ so that for all $\xi\in\Xi$, $E(\xi)$ have comparable cardinality. We may denote $\Xi\sim p^{s_{1}},\#E(\xi)\sim p^{s_{2}}$ for two dyadic numbers $p^{s_{1}},p^{s_{2}}\geq 1$. We can also assume $p^{s_{1}}p^{s_{2}}\gtrapprox\#E$. Hence, assuming (53) holds for $n^{\prime}=k+1$, we have
@@
However, we still need the condition $p^{s_{1}}\gtrsim p^{s}$. This is handled by the same trick as in the proof of Lemma 4.5. We will perform an algorithm to find a sequence $E=E_{1}\supset E_{2}\supset\cdots\supset E_{m}$, and disjoint subsets $\Xi_{1},\Xi_{2}\cdots,\Xi_{m}$ of $\mathbb{F}_{p}^{k}\times\mathbf{0}$. To start with, let $E_{1}=E$. By pigeonholing, we can find $\Xi_{1}\subset\mathbb{F}_{p}^{k}\times\mathbf{0}$, such that $\Xi_{1}\sim p^{u_{1}}$, $E(\xi)\sim p^{v_{1}}$ for each $\xi\in\Xi$, and
@@
Suppose we have defined $E_{i},\Xi_{i}$ for $1\leq i\leq j-1$. If $\sum_{i=1}^{j-1}p^{u_{i}}\ll p^{s}$, then we define
@@
Since each $V\in\mathcal{V}$ intersects $\xi+(\mathbf{0}^{k}\times\mathbb{F}_{p})$ (and hence $E(\xi)$) at a single point, we see that $E_{j}$ is still a $(s,t_{2};2,1)$-set. By the same pigeonholing, we find $\Xi_{j}\subset(\mathbb{F}_{p}^{k}\times\mathbf{0})\setminus\bigcup_{1\leq i\leq j-1}\Xi_{i}$, so that $\Xi_{j}\sim p^{u_{j}}$, $E(\xi)\sim p^{v_{j}}$ for each $\xi\in\Xi_{j}$, and
@@
If $\sum_{i=1}^{j-1}p^{u_{i}}\gtrsim p^{s}$, then we stop and let m=j1m=j-1.
The algorithm will stop in finite steps. Denote $\Xi:=\bigcup_{1\leq i\leq m}\Xi_{i}$. Let $p^{s_{1}},p^{s_{2}}$ be dyadic numbers so that $p^{s_{2}}=\min_{1\leq i\leq m}p^{v_{i}}$, $p^{s_{1}}\sim\#\Xi$. For each $\xi\in\Xi$, let $Y_{\xi}$ be a subset of $E(\xi)$ with $Y_{\xi}\sim p^{s_{2}}$. This is doable, since $E(\xi)\gtrsim p^{s_{2}}$ if $\xi\in\Xi$. One can check